OpenStudy (anonymous):

lim (1-x)^(1/8x) as x-->0

OpenStudy (anonymous):

draw it out for us. Your 1/8x looks dodgy. I don't know if the x is in the denominator or numerator.

OpenStudy (anonymous):

if you don't know how touse brackets efficiently so that people can see exactly what your equation/expression looks like, always go to Plan B and draw it out for us.

OpenStudy (anonymous):

to use*

OpenStudy (anonymous):

ive tried 1...says its wrong

OpenStudy (sirm3d):

do you recognize this? \[\Large \lim_{x\rightarrow 0}\left(1+x\right)^{\frac{1}{x}}\]

OpenStudy (anonymous):

a little

OpenStudy (sirm3d):

do you remember what the limit is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1/e?

OpenStudy (sirm3d):

\[\Large {\lim_{x\rightarrow 0} (1+x)^{\frac{1}{x}}=e\\\lim_{x\rightarrow 0} (1-x)^{\frac{1}{x}}=\frac{1}{e}}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

its 1/e^(1/8)

OpenStudy (sirm3d):

oh. great. saves me a lot of typing. =)